OpenStudy (anonymous):

How to solve? Find the Taylor series expansion of f(x)=(x-1)e^x about x=1.

OpenStudy (anonymous):

What am I supposed to do first? Thanks! :)

OpenStudy (anonymous):

(x-1) is already polynomial so leave it as it is

OpenStudy (anonymous):

Can you draw it please?

OpenStudy (anonymous):

draw what?

OpenStudy (anonymous):

in your text book , they might give you taylor expansion for e^x

OpenStudy (anonymous):

steps to solve this.

OpenStudy (anonymous):

We don't have books. :)) :(

OpenStudy (anonymous):

THANKS @imranmeah91 :)

OpenStudy (anonymous):

@imranmeah91 what does "about x=1" mean? Were you able to use that?

OpenStudy (anonymous):

it mean centered at x=1 =a so we have to use expanded form f(a)+f'(a)/1! (x-a)+ f''(a)/2! (x-a)^2 plug in 1 f(x)=e^x f(1)= e^1 e+ e (x-1)+ e(x-1)^2/2!+e(x-1)^2/3!

OpenStudy (anonymous):

so e^x centered at x=1 is \[e^x=\sum _{n=0}^{\infty } \frac{e(x-1)^n}{n!}\]

OpenStudy (anonymous):

multiply that by (x-1) \[e^x=\sum _{n=0}^{\infty } \frac{e(x-1)^{n+1}}{n!}\]